# What’s Wrong with Your Code Generated by Large Language Models? An Extensive Study

### v1

@@ 4.1. Taxonomy of Code Generation Bugs
4.1.4. Type C: Functional Bug.
Functional bugs refer to the bug in the program that causes it to behave incorrectly or not as intended according to its functional requirements (i.e., the code runs successfully but fails to pass all unit tests). Based on the taxonomy, there are four secondary functional bugs: Misunderstanding and Logic Error, Hallucination, Input/Output Format Error, and Minors.
C.1 Misunderstanding and Logic Error. Code generation tasks involve algorithmic problems where LLMs must extract information from natural languages and apply their knowledge to understand the requirements and establish correct logic. However, when faced with complex natural language descriptions, models often struggle to fully comprehend concepts, reference relationships, and conditional branches. For example, LLMs may incorrectly interpret integer concatenation as numerical addition. As shown in the code below, LLMs incorrectly interpret integer concatenation as numerical addition. What’s more, even if LLMs fully understand the problem description, converting this knowledge into correct logic remains challenging.
@@
C.2 Hallucination. Hallucination refers to instances where the LLM generates code that is syntactically plausible but factually incorrect or semantically meaningless. As shown in the below code, the code generated by the LLM does not align with the problem requirements at all.
@@
----- Canonical Solution -----
def is_Diff(n):
return n % 11 == 0
C.3 Input/Output Format Error. Different from bug type B.4 (i.e., incorrect argument), input/output format error refers to the incorrect order of inputs and outputs as well as the incorrect precision of the output data. As shown in the code below, the LLM incorrectly converts a float output to an integer.
@@
if human_years <= 2:
dog_years = human_years * 10.5
else:
dog_years = 21 + (human_years - 2) * 4
return int(dog_years) # Incorrect output format
@@
C.4 Minors. Minors in functional bugs include Incorrect Initialization, Sub-optimal Code, and Infinite Loop. Incorrect initialization indicates the code logic is correct, but incorrect initialization values for some variables prevent the code from passing unit tests. As shown in the code below, the variable max_sum is incorrectly initialized to 0. Sub-optimal code refers to instances where the LLM generates code using sub-optimal algorithms (e.g., greedy algorithms) to solve the problem, resulting in code that can only pass some unit tests but not all. Infinite loop refers to the code failing to meet the loop exit conditions under certain inputs, causing the code to run indefinitely.
Figure 4. Distribution of misunderstanding and logic error.
Write a function that returns the list in a list of lists whose sum of elements is the highest.
assert max_sum_list([[1,2,3], [4,5,6], [10,11,12], [7,8,9]])==[10, 11, 12]
------ Incorrect Code ------
def max_sum_list(lists):
max_sum = 0 # Incorrect initialization
max_list = None
for lst in lists:
current_sum = sum(lst)
if current_sum > max_sum:
max_sum = current_sum
max_list = lst
return max_list
----- Canonical Solution -----
def max_sum_list(lists):
return max(lists, key=sum)
4.2. Bug Analysis
We further analyzed the distribution of each primary and secondary bug type, as shown in Table 3. All data in the table represent the proportion of each bug type within the entire benchmark. Among the three primary bug types, syntax errors constitute the smallest proportion, whereas functional bugs are the most prevalent. As dataset complexity increases, the incidence of functional bugs also rises. Most LLMs have a syntax bug proportion of less than 10%, whereas DeepseekCoder, LLama-3, Phi-3, and GPT-3.5 have a functional bug proportion exceeding 50% on the APPS+. This suggests that current LLMs have effectively mitigated syntax bugs in code generation. Future efforts should prioritize enhancing the LLMs’ comprehension capabilities.
Finding 5: Among all LLMs and benchmarks, functional bugs constitute the highest proportion, whereas syntax bugs make up the lowest. Furthermore, as the complexity of the benchmark increases, the proportion of functional bugs also rises.
@@ 4.2. Bug Analysis
Among all secondary bug types, misunderstanding and logic errors make up the highest proportion, which is consistent across all LLMs and benchmarks. Furthermore, the more complex the benchmark, the more pronounced the LLMs’ shortcomings in understanding capabilities become. We further subdivided this bug type into six labels: Test-case-driven Code Generation, Missing Checks for Corner Cases, Reference Relationship Misunderstanding, Incorrect Conditional Branches, Specific Conception Misunderstanding, and Residual Logic Misunderstanding, as shown in Figure 4. Test-case-driven code generation indicates that to the LLM generating code solely based on the input and output of the test cases provided in the problem description, ignoring the actual functionality. Missing checks for corner cases indicates LLMs generate functionally correct code but fail to handle edge cases. Reference relationship misunderstanding indicates LLMs misinterpret reference relationships such as numerical relationships and order of operations. Incorrect conditional branches refers to the LLM only implementing or correctly handling some conditional branches. Specific conception misunderstanding refers to the LLM incorrectly understanding specific concepts defined in the problem description. The five labels above indicate that LLMs generally comprehend problem descriptions but make specific understanding errors. Residual logic misunderstanding refers to any remaining understanding errors not covered by the aforementioned five categories, as well as situations where the understanding is correct but the LLM fails to construct a correctly logical implementation of the code. The last label indicates that LLMs misunderstand the whole problem description.
From Figure 4, it can be observed that the missing checks for corner cases are substantial, accounting for 54.4% in HumanEval+ and 30.1% in MBPP+. Meanwhile, residual logic misunderstanding is most prevalent in the APPS+, where it reaches 69.1%, with a relatively high proportion in HumanEval+ and MBPP+. This indicates that LLMs find it more challenging to understand problem descriptions in the more complex APPS+. In contrast, their performances improve on HumanEval+, since they can grasp more problem details effectively.
Finding 6: Among all secondary bug types in LLMs and benchmarks, misunderstanding and logic errors are the most prevalent, and the distribution of the six sub-categories highlights LLMs’ weaknesses in comprehensive problem understanding.
Additionally, we found that as problem descriptions in benchmarks lengthen and become more complex, LLMs struggle to meet input-output format requirements and are more likely to generate incomplete syntax structures. This is because longer and more complex problem descriptions make it harder for the LLMs to remember details (Dou et al., 2024; Le et al., 2022), and the length of the generated code can lead to bugs like unmatched parentheses. This indicates that despite the large context windows of current LLMs (as shown in Figure 1), their ability to handle long texts is still insufficient, leading them to easily forget both the problem details and the content of the code already generated.
Finding 7: Despite the large context windows of current LLMs, their capacity to handle lengthy and complex problems is still limited. They struggle to retain all problem details, and the extended code generation results in more syntactical errors like unmatched parentheses.
@@
API misuse accounts for the highest proportion of runtime bugs across most benchmarks and large language models (LLMs). There are three primary errors associated with API misuse: AttributeError (20.9%), TypeError (50%), and ValueError (26.9%). An AttributeError occurs when an invalid reference is made to a missing API attribute. A TypeError arises when an API is applied to an object of an incorrect type, while a ValueError occurs when an argument of the correct type but inappropriate value is provided Thus, LLMs must accurately infer multiple factors simultaneously when invoking APIs, including the attributes of the caller, the types of API arguments, and the appropriate range of argument values. Therefore, future work can focus on improving the LLM’s inference of variable attributes, types and values to prevent API misuse.
@@
Finding 8: API misuse represents the most common runtime bug in benchmarks and LLMs, necessitating accurate inference of caller attributes, argument types, and value ranges.
Timeout errors indicate that while the LLM-generated code is correct, the algorithm’s implementation is often suboptimal, typically with high time complexity, preventing the program from finishing within the time limit. For more complex algorithmic problems, LLMs tend to use loops or deep searches without optimized exit conditions. Additionally, the frequent use of recursive algorithms significantly increases the time complexity.
1def fib(n: int):
2 if n <= 1:
3 return n
4 else:
5 return fib(n-1) + fib(n-2)
1def fib(n: int):
2 if n == 0: return 0
3 if n <= 2: return 1
4 a, b = 1, 1
5 for _ in range(3, n + 1):
6 a, b, = b, a + b
7 return b
Finding 9: For complex problems, LLMs often fail to generate optimal algorithms, frequently using loops or deep searches without optimized exit conditions or recursive algorithms, leading to timeout errors.
Incorrect argument bug type is uncommon, but in the APPS+, 37.7% of StarCoder-2’s code contains this issue. Specifically, while StarCoder-2 understands input formats, it fails to generate redundant arguments as required by the problem statement. The redundant argument refers to argument that can be inferred from other parameters. As illustrated in the final code example in Section 4.1.3, the incorrect code does not follow input requirements and ignores the first argument, which can be inferred from the length of the second argument.
Finding 10: Incorrect argument is rare in most LLMs while StarCoder-2 usually fails to generate redundant arguments required by the problem statement.
Figure 5. The process of constructing RWPB.
5. Code Generation in Real-World Projects
Constructing a real-world code generation benchmark is crucial for evaluating the effectiveness of LLMs in generating code applicable to real-world projects. Moreover, the distribution of bugs in the incorrect code may vary between the existing popular benchmarks and the real-world benchmark, emphasizing the importance of the real-world benchmark in understanding and improving LLM performance in real-world scenarios. In this section, we first introduce the process of constructing our real-world benchmark RWPB. Subsequently, we evaluate widely used LLMs’ performance in code generation on RWPB and discuss the challenges and limitations inherent in these LLMs for real-world projects.
5.1. Benchmark Construction
@@
Our constructed benchmark RWPB contains 140 real-world programming requirements collected from Github. Each requirement includes a function signature, docstring (i.e., descriptions of the function’s functionality and input-output), function body, and unit test cases. As shown in Figure 5, the constructing process included five steps:
@@
(1) Repository Selection. Currently, commonly used benchmarks are constructed from programming questions, which are disconnected from real-world program development. In addition, other real-world benchmarks are sourced from widely used real-world repositories and libraries. These benchmarks suffer from data contamination and leakage due to significant overlap with the datasets used for training LLMs (Chen et al., 2021; Riddell et al., 2024; Roberts et al., 2023; Golchin and Surdeanu, 2023). To address the shortcomings of current benchmarks, we collected Python repositories established in 2024 on GitHub. Specifically, we focused on repositories created before April 15, 2024, and filtered out those with licenses that restricted open-source or commercial uses. We then ranked the retained repositories in descending order of the number of stars, and ultimately obtained 600 practical repositories.
@@
(2) Function Filtering. We extracted functions from these projects and wrote automated scripts to filter out empty or initialization functions. We also removed functions that incorporated private APIs and dependencies, exceeded 100 lines of code, and contained many constants like strings and numbers. We found that most of the retaining functions existed in the utility files of projects. In this stage, we obtained 2,307 candidate functions.
@@
(3) Function De-Duplication. Developers frequently construct their projects using existing popular tools and code repositories. However, these codes may result in data leakage since current LLMs are trained on a large fraction of GitHub and libraries. To solve this challenge, we utilized the MinHash algorithm (Broder, 1997) to ensure the absence of duplication between functions, as well as between these functions and code bases, including open-source bases (i.e., RedPajama Data v2, Stack v1 (Kocetkov et al., [n. d.]) and v2 (Lozhkov et al., 2024), RefinedWeb (Penedo et al., 2023), and Pile (Gao et al., 2020)), and our internal code training corpora. This method significantly reduces the potential for data leakage in our benchmark. Finally, we retained 316 candidate functions.
@@
(4) Human Annotation. After removing potentially leaked functions, we annotated the docstring for these functions. We engaged nine experts to manually annotate each function’s docstring, including descriptions of its functionality and input-output. The annotated docstring follows three criteria: naturalness, accuracy, and comprehensiveness. Specifically, the docstring should follow the human writing style, making it naturally readable from a developer’s perspective. Additionally, it should comprehensively describe the function’s functionality and input-output parameters, such as containing input-output types and covering all conditional branches. Finally, the docstring should be clear, precise and concise, avoiding any redundant and irrelevant information.
Each docstring involved a dual annotation process and three annotators: one expert manually wrote a docstring and the others provided a meticulous double-check. If the two experts responsible for the double-check provided differing results, the drafted docstring would be manually refined until a consensus was reached. During this process, experts were allowed to utilize GPT-4 to assist with annotation tasks. Specifically, the first expert was permitted to use GPT-4 to help summarize the functionality, while other experts were allowed to use it to aid in the check process. Experts also eliminated functions that exhibited a conglomeration of functionalities and were difficult to summarize exactly and thoroughly. Finally, this stage resulted in the retention of 140 functions, each containing an accurate and comprehensive docstring.
@@
(5) Unit Tests Synthesis. We engaged six experts to create unit test cases for each function. The generation of test cases is based on the following key criteria: Each function contains at least three test cases; these test cases are ensured to cover the entire code of the function body; and they comprise regular and boundary cases. In addition, as in Stage Four, we implemented a double-check synthesis process. Specifically, for each function, one expert undertook the task of generating test cases, while subsequently, two other experts validated the synthesized cases and checked whether they could cover the full canonical solution of this function. At the end of the stage, on average, each function contained 4.9 test cases.
@@
Table 4. Types of bugs introduced during code generation by seven popular LLMs on our constructed real-world benchmark RWPB. All values are in %.
@@
StarCoder2 DeepSeekCoder Llama3 Phi3 GPT4 GPT3.5 Claude3
PASS 31.4 32.9 22.9 22.1 44.3 34.3 45.7
A.1 0.0 0.0 0.7 0.0 0.7 0.7 0.0
A.2 2.1 0.0 0.7 2.9 2.1 2.1 0.0
A.3 4.3 2.9 2.9 2.1 2.9 4.3 1.4
A 6.4 2.9 4.3 5.0 5.7 7.1 1.4
B.1 7.9 7.1 10.7 10.0 2.9 5.0 2.9
B.2 1.4 0.7 0.7 5.7 1.4 2.1 7.9
B.3 5.0 4.3 3.6 2.1 1.4 3.6 2.9
B.4 12.9 19.3 19.3 16.4 7.1 12.1 7.1
B.5 0.0 1.4 3.6 2.9 0.0 1.4 1.4
B 27.1 32.9 37.9 37.1 12.9 24.3 22.1
C.1 27.1 21.4 24.3 28.6 26.4 25.7 25.0
C.2 6.4 5.7 7.1 4.3 5.0 6.4 0.0
C.3 1.4 2.1 0.7 0.7 2.1 1.4 3.6
C.4 0.0 2.1 2.9 2.1 3.6 0.7 2.1
C 35.0 31.4 35.0 35.7 37.1 34.3 30.7
5.2. Effectiveness in RWPB and Bugs Analysis
5.2.1. Effectiveness.
As shown in Table 4, among closed-source LLMs, Claude-3 has the highest pass rate at 45.7%, followed by GPT-4 at 44.3%. For open-source LLMs, StarCoder-2 and DeepSeekCoder have the highest pass rates at 31.4% and 32.9%, respectively, while Llama-3 and Phi-3 have lower pass rates at 22.9% and 22.1%, respectively. Open-source LLMs have a slightly lower overall pass rate compared to closed-source LLMs, particularly performing worse in runtime bugs. This indicates that closed-source LLMs can effectively avoid runtime bugs in generated code, which also means they are more likely to expose functional bugs due to the annotation priority. The differences among LLMs in functional bugs are minimal. We further investigated the overlap of functional bugs for each LLM. Among the seven LLMs, 100 problems functional bugs and out of these, only seven problems have functional bugs reported by every model. This indicates that, for real-world projects, different LLMs have varying areas of focus in real-world scenarios. This variability is likely due to differences in the source of training data, LLM size, training methods, and hyper-parameter details, which result in varying levels of understanding across models.
Finding 11: Closed-source LLMs outperform open-source LLMs in code generation quality for real-world projects, particularly excelling in reducing syntax bugs and runtime bugs. For functional bugs, different LLMs perform similarly, with each model encountering errors on different problems.
5.2.2. Bug Analysis.

### v2

@@ 4.1.1 Methodology
4.1.4 Type C: Functional Bug.
Functional bug refers to the bug in the program that causes it to behave incorrectly or not as intended according to its functional requirements (i.e., code runs successfully but fails to pass all unit tests). Based on the taxonomy, there are four secondary functional bugs: Misunderstanding and Logic Error, Hallucination, Input/Output Format Error, and Minors.
C.1 Misunderstanding and Logic Error. Code generation tasks involve algorithmic problems where LLMs must extract information from natural languages and apply their knowledge to understand the requirements and establish correct logic. However, when faced with complex natural language descriptions, models often struggle to fully comprehend concepts, reference relationships, and conditional branches. As shown in the code below, LLMs incorrectly interpret integer concatenation as numerical addition. What is more, even if LLMs fully understand the problem description, converting this knowledge into correct logic remains challenging.
@@
C.2 Hallucination. Hallucination refers to instances where the LLM generates code that is syntactically plausible but functionally incorrect or semantically meaningless. As shown in the code below, the code generated by the LLM does not align with the problem requirements.
@@
C.3 Input/Output Format Error. Different from bug type B.4 (i.e., incorrect argument), input/output format error refers to the incorrect order of inputs/outputs and the incorrect precision of the output data. As shown in the code below, the LLM incorrectly converts a float output to an integer.
@@
return int(dog_years) # Incorrect format
@@
C.4 Minors. Minors in functional bugs include Incorrect Initialization, Sub-optimal Code, and Infinite Loop. Incorrect initialization indicates the code logic is correct, but incorrect initialization values for some variables prevent the code from passing all unit tests. Sub-optimal code refers to instances where LLMs generate code using sub-optimal algorithms (e.g., greedy algorithms) to solve the problem, resulting in code that can only pass partial unit tests. Infinite loop refers to the generated code failing to meet the loop exit conditions under certain inputs, causing the code to run indefinitely.
Figure 4: Distribution of misunderstanding and logic error.
4.2 Bug Analysis
Bug Distribution. We further analyzed the distribution of primary and secondary bug types, as shown in Table 3 and Table 4. All results in the table represent the proportion of each bug type within the entire benchmark. Among the three primary bug types, syntax errors constitute the smallest proportion, whereas functional bugs are the most prevalent. This pattern holds consistently across both open-source and closed-source models, with a notable exception: DeepSeek-R1 achieves remarkably low functional bug rates (6.1% on HumanEval+, 12.7% on MBPP+, and 25.0% on APPS+), significantly outperforming other models, including GPT-4 (13.4%, 16.5%, and 39.3%, respectively) and Claude-3 (19.5%, 19.8%, and 45.7%, respectively). The superior performance of DeepSeek-R1 becomes more pronounced as dataset complexity increases. While most models exhibit a sharp increase in functional bugs on APPS+, with DeepSeekCoder, LLaMA-3, Phi-3, and GPT-3.5 each exceeding 50% functional bug rates, DeepSeek-R1 maintains the lowest rate at 25%, representing a 36% relative improvement over GPT-4 and a 45% improvement over Claude-3. This suggests that DeepSeek-R1 possesses enhanced semantic understanding capabilities that remain robust even when faced with complex problem descriptions. Examining runtime bugs, DeepSeek-R1 demonstrates competitive performance with rates of 0.6%, 5.6%, and 4.7% across the three benchmarks. Particularly noteworthy is its consistent avoidance of certain bug types: it produces zero hallucinations across all benchmarks, minimal incorrect arguments, and maintains low API misuse rates. This pattern indicates systematic improvements in code reliability and adherence to programming constraints.
Figure 5: The process of constructing RWPB.
@@ 4.2 Bug Analysis
Among all secondary bug types, misunderstanding and logic errors form the largest proportion, which is consistent across all LLMs and benchmarks. However, DeepSeek-R1 achieves the lowest rates in this category across all closed-source models: 5.5% (HumanEval+), 12.4% (MBPP+), and 21.5% (APPS+), compared to GPT-4’s 12.8%, 12.0%, and 31.3%. This represents a 57% reduction in logic errors on HumanEval+ and a 31% reduction on APPS+ compared to GPT-4, suggesting that DeepSeek-R1’s architecture or training methodology specifically addresses the comprehension challenges that plague other models. We further subdivided the misunderstanding and logic error category into six labels: Test-case-driven Code Generation, Missing Checks for Corner Cases, Reference Relationship Misunderstanding, Incorrect Conditional Branches, Specific Conception Misunderstanding, and Residual Logic Misunderstanding, as shown in Figure 4. Test-case-driven code generation indicates LLMs generate code solely based on the test cases in the problem description, ignoring the actual functionality. Missing checks for corner cases indicate LLMs generate functionally correct code but fail to handle edge cases. Reference relationship misunderstanding indicates LLMs misinterpret reference relationships, such as numerical relationships and order of operations. Incorrect conditional branches indicate that LLMs only correctly handle some conditional branches. Specific conception misunderstanding indicates LLMs incorrectly understand specific concepts defined in the problem description. Residual logic misunderstanding refers to any remaining cases not covered by the previous labels, including instances where the understanding is correct, but the LLM fails to implement the logic correctly.
From Figure 4, missing checks for corner cases account for a large portion of errors in HumanEval+ (54.4%) and MBPP+ (30.1%). Meanwhile, residual logic misunderstanding is the most prevalent in APPS+, where it reaches 69.1%, with a relatively high proportion in HumanEval+ and MBPP+. This indicates that it is more challenging for LLMs to understand problem descriptions in more complex APPS+. The exceptional performance of DeepSeek-R1, particularly its low rates of logic errors and complete absence of hallucinations, suggests that it may employ advanced reasoning mechanisms or enhanced training strategies that enable more accurate interpretation of complex problem specifications. This positions DeepSeek-R1 as a particularly promising model for code translation tasks requiring deep semantic understanding, setting a new benchmark for future LLM development in software engineering applications.
Finding 5: Functional bugs account for the highest proportion, while syntax bugs are the least common. As the complexity of the benchmark increases, the proportion of functional bugs also rises. Among secondary bug types, misunderstanding and logic errors are the most prevalent. Implication: Enhancing an LLM’s ability to correctly interpret problem statements, especially corner cases and nuanced logic, represents a critical next step in code generation research. Researchers can also introduce critique techniques [63] to understand or verify each code segment to improve the correctness.
@@
API Misuse. API misuse accounts for the highest proportion of runtime bugs across most benchmarks and large language models (LLMs). There are three primary errors associated with API misuse: AttributeError (20.9%), TypeError (50%), and ValueError (26.9%). An AttributeError occurs when an invalid reference is made to a missing API attribute. A TypeError arises when an API is applied to an object of an incorrect type, while a ValueError occurs when an argument of the correct type but inappropriate value is provided. Thus, LLMs must accurately infer multiple factors simultaneously when invoking APIs, including the attributes of the caller, the types of API arguments, and the appropriate range of argument values. Therefore, future work can focus on improving the LLM’s inference of variable attributes, types and values to prevent API misuse.
@@
Finding 6: API misuse represents the most common runtime bugs, necessitating accurate inference of caller attributes, argument types, and value ranges. Implication: Future research should enhance LLMs’ capacity to infer and validate all components involved in API calls, i.e., object attributes, argument types, and permissible value ranges, to reduce runtime bugs and strengthen overall code reliability.
Timeout Error. A timeout error indicates that the LLM-generated code is functionally correct yet fails to complete its execution within the specified time limit, primarily due to suboptimal algorithmic choices. In complex tasks, models frequently implement nested loops or conduct deep searches without sufficient optimization or effective exit conditions, leading to excessive time complexity. The repeated use of recursive algorithms, often without appropriate bounding logic, further increases the computational cost.
Finding 7: Timeout errors are mainly caused by inefficient or unoptimized implementations, especially in complex problems where loops and recursive searches significantly increase time complexity. Implication: Future research should focus on LLM algorithmic optimization and complexity awareness to avoid surpassing time limits.
5 Code Generation in Real-World Projects
Constructing a real-world code generation benchmark is essential for assessing LLMs’ effectiveness in generating practical code. The distribution of bugs in real-world code may differ from that in popular benchmarks, highlighting the need for a real-world benchmark to better understand and enhance LLMs’ performance. This section introduces the RWPB benchmark, evaluates the performance of widely used LLMs on it, and discusses the associated challenges and limitations.
5.1 Benchmark Construction
@@
Our constructed benchmark RWPB contains 140 real-world programming requirements collected from GitHub. Each requirement includes a function signature, docstring (i.e., the problem description, arguments, and returns), function body, and unit test cases. As shown in Figure 5, the construction process included the following five main steps:
@@
(1) Repository Selection. Currently, commonly used benchmarks are constructed from programming questions, which are disconnected from real-world program development. In addition, other real-world benchmarks are sourced from widely used real-world repositories and libraries. These benchmarks suffer from data contamination and leakage due to significant overlap with the datasets used for training LLMs [31, 37, 38, 35]. To address the shortcomings of current benchmarks, we collected Python repositories established in 2024 on GitHub. Specifically, we focused on repositories created before April 15, 2024, and filtered out those with licenses that restricted open-source or commercial uses. We then ranked the retained repositories in descending order of the number of stars, and ultimately obtained 600 practical repositories.
@@
(2) Function Filtering. We extracted functions from these projects and filtered out empty or initialization functions. Following the filter approach of HumanEval [31], we also removed functions that incorporated private APIs and dependencies (i.e., that are not intended to be publicly accessible or used by developers), exceeded 100 lines of code, and contained many constants like strings and numbers. In this stage, we obtained 2,307 candidate functions.
@@
(3) Function De-Duplication. Developers frequently construct their projects using existing popular tools and code repositories. However, such code may result in data leakage since current LLMs are trained on a large fraction of GitHub and libraries. To solve this challenge, we conduct de-duplication between our collected functions and widely used open-source code bases (i.e., RedPajama Data v2, Stack v1 [64], Stack v2 [18], RefinedWeb [65], and the Pile [66]) by utilizing the MinHash algorithm [67]. This method significantly reduces the potential for data leakage in our benchmark. Finally, we retained 316 candidate functions.
@@
(4) Human Annotation. After eliminating potentially leaked functions, we annotated the docstring for these functions. We engaged nine experts to manually annotate each function’s docstring, including functionality descriptions, arguments, and returns. The annotated docstring follows three criteria: naturalness, accuracy, and clarity. Specifically, firstly, the docstring should follow the human writing style, making it naturally readable from a developer’s perspective. Secondly, the description should accurately describe the function’s functionality. The arguments and return values also should be accurate, such as the type and name should be identical to the code. Finally, the docstring should be clear, precise, and concise, avoiding any redundant and irrelevant information. Each docstring involved a dual annotation process and three annotators: one expert manually wrote a docstring, and the others provided a meticulous double-check. If the two experts responsible for the double-check provided differing results, the drafted docstring would be manually refined until a consensus was reached. Experts also eliminated functions that exhibited a conglomeration of functionalities and were difficult to summarize exactly and thoroughly. Finally, this stage retained 140 functions, each containing an accurate and comprehensive docstring.
@@
(5) Unit Tests Synthesis. We engaged six experts to write unit test cases for each function. The generation of test cases is based on the following key criteria: Each function contains at least three test cases; these test cases are ensured to cover the entire code of the function body; and they comprise regular and boundary cases. In addition, similar to Stage Four, we implemented a double-check synthesis process. For each function, one expert was responsible for writing test cases. Subsequently, two experts review all samples to ensure the coverage of the unit tests is 100%.
@@
Table 5: Types of bugs introduced during code generation by nine popular LLMs on our constructed real-world benchmark RWPB. DSV denotes DeepSeek-V3. DSR denotes DeepSeek-R1. All values are in %.
@@
StarCoder2 DeepSeekCoder Llama3 Phi3 DSV DSR GPT4 GPT3.5 Claude3
PASS 31.4 32.9 22.9 22.1 75.7 77.9 44.3 34.3 45.7
A.1 0.0 0.0 0.7 0.0 0.0 0.0 0.7 0.7 0.0
A.2 2.1 0.0 0.7 2.9 0.0 0.0 2.1 2.1 0.0
A.3 4.3 2.9 2.9 2.1 1.4 2.1 2.9 4.3 1.4
A 6.4 2.9 4.3 5.0 1.4 2.1 5.7 7.1 1.4
B.1 7.9 7.1 10.7 10.0 7.9 3.6 2.9 5.0 2.9
B.2 1.4 0.7 0.7 5.7 0.0 0.0 1.4 2.1 7.9
B.3 5.0 4.3 3.6 2.1 0.7 0.7 1.4 3.6 2.9
B.4 12.9 19.3 19.3 16.4 0.0 1.4 7.1 12.1 7.1
B.5 0.0 1.4 3.6 2.9 0.0 0.0 0.0 1.4 1.4
B 27.1 32.9 37.9 37.1 8.6 5.7 12.9 24.3 22.1
C.1 27.1 21.4 24.3 28.6 12.9 11.4 26.4 25.7 25.0
C.2 6.4 5.7 7.1 4.3 0.0 0.0 5.0 6.4 0.0
C.3 1.4 2.1 0.7 0.7 1.4 2.1 2.1 1.4 3.6
C.4 0.0 2.1 2.9 2.1 0.0 0.7 3.6 0.7 2.1
C 35.0 31.4 35.0 35.7 14.3 14.3 37.1 34.3 30.7
5.2 Effectiveness in RWPB and Bugs Analysis
Effectiveness. As shown in Table 5, the evaluated LLMs demonstrate significant performance variation. DeepSeek-R1 and DeepSeek-V3 achieve the highest passing rates at 77.9% and 75.7%, respectively, substantially outperforming all other models. Among the remaining closed-source LLMs, Claude-3 reaches 45.7% and GPT-4 achieves 44.3%, while GPT-3.5 lags behind at 34.3%. Open-source models show notably lower performance, with StarCoder-2 and DeepSeekCoder achieving passing rates of 31.4% and 32.9%, respectively, while Llama-3 and Phi-3 trail at 22.9% and 22.1%.
A key observation emerges from analyzing the distribution of bug types across models. DeepSeek-R1 and DeepSeek-V3 excel particularly in avoiding runtime bugs, with occurrence rates of only 5.7% and 8.6% respectively, significantly lower than the 12.9%-37.9% range observed in other models. This superior performance suggests these models have developed a more robust understanding of execution semantics. Notably, both DeepSeek models demonstrate exceptional capability in handling Type B.4 bugs, which typically involve complex control flow issues. When excluding the high-performing DeepSeek models, a clear pattern emerges: traditional open-source models exhibit higher rates of runtime bugs (27.1%-37.9%) compared to their closed-source counterparts (12.9%-24.3%), indicating that closed-source LLMs generally possess better runtime semantic understanding.
In contrast, the performance gap narrows considerably for functional bugs. Most models cluster within a similar range of 30.7% to 37.1%, though DeepSeek-V3 and DeepSeek-R1 maintain their advantage with only 14.3% occurrence. Further investigation into the overlap of functional bugs across LLMs revealed limited commonality, suggesting that different models possess distinct areas of expertise in real-world scenarios. This variability likely stems from differences in training data sources, model architectures, training methodologies, and hyperparameter configurations, which collectively shape each model’s unique understanding and capabilities.
Finding 8: Closed-source LLMs outperform open-source LLMs in generated code quality for real-world projects, particularly excelling in reducing syntax and runtime bugs. Implication: Researchers should incorporate a diverse range of real projects into the training set to generate the generalization of LLMs on real-world benchmarks. The substantial performance gains observed in reasoning-enhanced models suggest that future research should explore hybrid training approaches that combine supervised learning on large-scale code corpora with reinforcement learning from compiler feedback, test execution results, and static analysis tools.